# The Cosmological CPT Theorem

### v2

@@ 1 Introduction
The reason we can obtain these stronger constraints, is that we are using the additional conditions that (i) we are considering the vacuum state only, and (ii) the dynamics of the theory come from a local Lagrangian that has CRT symmetry around every point. This combination of discrete symmetries with locality turns out to give results much more powerful than those which can be obtained by simply thinking of discrete symmetries as acting globally on correlators, without considering the aetiology of the correlators, i.e. the manner in which they arise from interactions, starting from a Bunch-Davies boundary condition.
@@ 1.2 Applications to Cosmology
[…]
@@
where the last term indicates that $\bf{CRT}$ can only fix the phase up to an overall $\pm$ sign. $\Delta$ is the conformal dimension of the $n$ operators $\overline{\phi}_{+}$ used to differentiate the boundary wavefunction $\overline{\Psi}[\overline{\phi}_{-}]$, which is written in the basis of the conjugate operators $\overline{\phi}_{-}$ with dimension $d-\Delta$ [^9] From a dS/CFT perspective, the objects with dimension $d-\Delta$ are the “sources” of the CFT partition function $Z_{\text{CFT}}=\overline{\Psi}$, while the ones with dimension $\Delta$ are the “operators”. But from a cosmology perspective, both are operators in the bulk Hilbert space, and we have to choose a basis. In cosmology, one normally chooses this basis by writing $\overline{\Psi}$ as a function of whichever field component falls off more slowly, which would imply that $\Delta\geq d/2$, but the phase formula above would still be valid if we make the opposite choice. We do not attempt in this paper to extend our results to heavy fields in the principal series, where $\Delta=d/2+i\nu$. Doing so would require making the unpleasant decision to either sacrifice self-adjointness, or conformal invariance of the boundary conditions. See e.g. [76, 77, 78, 79, 80, 81, 82, 83, 84, 85] for discussions regarding the principal series. This result holds for any boundary wavefunction coefficients that ae UV- and IR-finite (i.e. free from logarithmic divergences in $\eta$ or the UV cutoff $\epsilon$), as well as the coefficient in front of the leading $\log(\eta)$ divergence for any IR-divergent wavefunction coefficients, or the leading $\log(\epsilon)$ divergence for any UV-divergent diagram. [^10] There is also an extra correction for diagrams with spinor propagators, see footnote 43.
@@
This then tells us that massless scalar fields and gravitons have real boundary wavefunction coefficients to all loop order in even spacetime dimension provided they don’t diverge in the UV. When UV divergences are present one finds in dimensional regularisation (dim-reg) a generic complex phase proportional to the dim-reg parameter $\delta$, i.e. $\arg(\overline{\psi}^{(L)}_{n})\propto\pi\delta$. For tree-level, IR-finite and scale-invariant Feynman-Witten diagrams in cosmology, this was recently found independently [74, 75] where the authors proved this reality condition by direct analysis of the bulk time integrals. [^11] AT would like to express his gratitude to David Stefanyszyn, Xi Tong and Yuhang Zhu for various extensive and helpful discussions on reality in cosmology. The implications of this for cosmology in terms of no-go theorems for parity violations has previously been explored for the specific case of the tree-level scalar trispectrum in [86, 87], and will be treated for more general tree-level and loop diagrams in [88]. This is also consistent with what is found in flat space amplitudes, which can be shown to be real at tree-level, see e.g. [89]. [^12] AT thanks Carlos Duaso Pueyo for making him aware of this result in amplitudes.
@@ 1.3 Implications for dS/CFT
These constraints not only have implications for cosmology [88] but may also illuminate the role of Unitarity in de Sitter quantum gravity (dS/CFT), where previous attempts to identify the imprints of bulk unitary time evolution in the boundary theory had proven to be elusive (see e.g. [25, 26]). Given the close relationship between $\bf{CRT}$ and Unitarity, and the fact that $\bf{CRT}$ simplifies at the future boundary ${\cal I}^{+}$, our work gives some reason to think that bulk Unitarity corresponds to some equally natural restriction on the boundary CFT side. At the very least, if somebody claims to have an example of dS/CFT holography, the phase formula would give a simple way to rule out examples where the bulk theory violates Reflection Reality ($\bf{RR}$), without any need to do calculations that probe the bulk interior. Then, to ensure that the bulk theory is fully unitary, the remaining task is to check by hand that that the bulk Lagrangian has no negative norm fields. We expect that success at this task would not require any additional fine-tuning of continuous parameters, although it might require making good discrete choices. [^13] Of course, there are some other consistency conditions that dS/CFT should obey besides bulk Unitarity. This includes normalisability conditions on the wavefunction $\Psi$, which can be used to fix the sign of the real part of typical $\psi_{2}$ coefficients. A particularly strong constraint is the absence of bulk tachyons, which seems to require the absence of irrelevant single-trace operators $\Delta>d$ in the boundary field theory.
@@
  • For $d=2$ at tree-level, $c=i\frac{3\ell}{2G_{N}}$ (e.g. from applying Friedel’s analysis to dS [59]); while for the case of pure gravity at 1-loop order we have $c=i\frac{3\ell}{2G_{N}}+13+\mathcal{O}(G_{N})$ [^15] The renormalisation of Newton’s constant $G_{N}$ is scheme dependent, meaning its exact running or loop corrections depend on the choice of renormalisation procedure. In (A)dS3, when one reformulates the gravitational path integral (GPI) using a boundary Schwarzian-type description, there exists a localisation result (analogous to the Duistermaat–Heckman theorem) which implies that, in that specific scheme, the renormalisation of $G_{N}$ stops at one loop. However, this one-loop exactness is not universal and might not hold in other renormalisation schemes. Importantly, the physical quantity is the boundary central charge (which appears in the dual CFT), not the bare $G_{N}$ itself. Therefore, as is standard in QFT, one should treat $G_{N}$ as a scheme-dependent parameter that is adjusted (renormalised) to keep physical observables—like the central charge—fixed. [91]; [^16] We thank Jordan Cotler and Kristan Jensen for making us aware of their previous result for the 1-loop contribution to the central charge in $dS_{3}$ as well as upcoming work for loop calculations in $dS_{2}$ and $dS_{4}$. The more observant reader will notice that the real 111-loop contribution is positive in the convention chosen by the authors as predicted from the combination of the normalisability constraint and the phase formula.
@@ 2.2 Outline of CPT Theorem
Assumption 3 allows us to analytically continue to Euclidean signature, using the fact that $e^{-H\Delta\tau}$ is suppressed for positive Euclidean times $\Delta\tau>0$. Let us adopt coordinates in Euclidean signature where $\tau=it$ and $x_{I}$, $I=1\,...\,d$ are the spatial coordinates. We highlight here (as it will become important in the cosmological case) that the $\bf{R}$ in $\bf{CRT}$ reflects a single spatial coordinate, say $x_{1}\rightarrow-x_{1}$.
The $SO^{+}(d,1)$ Lorentz Invariance in $d+1$-dimensional Minkowski spacetime analytically continues to $SO(d+1)$ rotational invariance in Euclidean signature. On the other hand, Unitarity implies that Euclidean correlators satisfy Reflection Positivity, which we will define in the following section. However, it is not necessary to enforce either of these relationships in their entirety to ensure $\bf{CRT}$ symmetry. We do not require full $SO^{+}(d,1)$ Lorentz Invariance; we simply need the subgroup $SO^{+}(1,1)$ boost  [^18] Consider a Lorentz boost in the $t$-$x^{d}$ plane of $d+1$-dimensional Minkowski spacetime. The infinitesimal transformation of coordinates is given by $\begin{pmatrix}t^{\prime}\\ x^{D}{}^{\prime}\end{pmatrix}=\begin{pmatrix}\cosh\xi&-\sinh\xi\\ -\sinh\xi&\cosh\xi\end{pmatrix}\begin{pmatrix}t\\ x^{D}\end{pmatrix},$ (2.1) where $\xi$ is the rapidity parameter of the boost, and we have omitted dependence on the other $d-1$ spatial coordinates for simplicity. To Wick rotate to Euclidean signature, we set $t=-i\tau\qquad\text{and}\qquad\xi=-i\theta\,,$ (2.2) which leads to the transformation $\begin{pmatrix}-i\tau^{\prime}\\ x^{D}{}^{\prime}\end{pmatrix}=\begin{pmatrix}\cosh(-i\theta)&\sinh(-i\theta)\\ \sinh(-i\theta)&\cosh(-i\theta)\end{pmatrix}\begin{pmatrix}-i\tau\\ x^{D}\end{pmatrix}.$ (2.3) Using the identities $\cosh(-i\theta)=\cos\theta$ and $\sinh(-i\theta)=-i\sin\theta$, this becomes $\begin{pmatrix}\tau^{\prime}\\ x^{D}{}^{\prime}\end{pmatrix}=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\begin{pmatrix}\tau\\ x^{D}\end{pmatrix},$ (2.4) which is a clockwise rotation in the $\tau$-$x^{D}$ plane by angle $\theta$. Hence, under the Wick rotation $t=-i\tau$ and $\xi=-i\theta$, the non-compact Lorentz boosts $SO^{+}(1,1)$ in Minkowski space analytically continue to compact Euclidean rotations $SO(2)$ in the $\tau$-$x^{D}$ plane. (this will become important when studying the implications of the CPT theorem for cosmology). In fact, all that is required is a discrete, $\mathbb{Z}_{2}$ symmetry from each of them:
@@
  • A $180^{\circ}$ ($=\pi$) rotation in the $\tau$-$x$ plane corresponds to the transformation: [^20] Some authors refer to this transformation as “CPT”. The difference in convention goes back to the earliest papers on the CPT theorem, see [95] for a nice review. In the language of [95], we are defining $\bf{T}$ as a “Wigner time reversal” (which involves a complex conjugation) while other authors use “Schwinger time reversal” (which instead reverses the order of operators. For example, the recent paper by Susskind [96] and [97], as well as Harlow and Numasawa [98] on gauging “CRT”, is actually defining $\bf{CRT}$ as what we call the $180^{\circ}$ rotation, at least in their Euclidean signature construction (see also [99], where the authors find that in order to construct physically consistent bulk local states, it is necessary to take a CPT-invariant linear combination of the two primary states of the dual CFT). It is unclear whether it makes sense to gauge our Wignerian $\bf{CRT}$, as this would raise the question of what it means to have a non-trivial complex conjugation of amplitudes when going around a non-contractible cycle. (Perhaps it just means the whole amplitude vanishes.)
@@
For theories involving spinor fields, it is also necessary to consider projective representations of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. In this paragraph, let us assume for simplicity that the theory is fully rotationally covariant. Then a $360^{\circ}$ rotation gives a factor of $(-1)^{2s}$ (where $s$ is the spin); this is of course the same as the fermion parity $(-1)^{F}$ for fields that satisfy spin-statistics. Hence, the $180^{\circ}$ rotation $R_{\pi}$ has order 444 when acting on half-integer spin fields, whilst Unitarity requires that ${\bf RR}^{2}=+1$ and hence in a rotationally covariant theory we also need ${\bf{CRT}}^{2}=+1$. This is only possible if the group anticommutes in the sense that $[A,B]:=ABA^{-1}B^{-1}=-1$ when $A,B$ are any two distinct elements of $\bf{RR}$, $\bf{CRT}$, ${\bf R}_{\pi}$}. One thus finds that the appropriate double cover of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ for spinors is the dihedral group $D_{4}$. [^22] With the addition of one additional magical fact that Unitarity requires ${\bf CRT}^{2}=(-1)^{2s}(-1)^{F}$, which requires a close inspection of the transformation properties of 2-pt correlators to prove, the considerations in this paragraph provide all the necessary ingredients to prove the spin-statistics theorem. But spin-statistics requires at least covariance under $90^{\circ}$ rotations, not just $180^{\circ}$ rotations, so it is less robust than the CPT theorem. We will focus on fields with integer spin representations and not consider half-integer spin fields in detail this paper, but a similar formalism should also provide a much easier approach to obtain a Unitarity constraint for correlators involving spinors in flat space.
@@ 2.3 Reflection Positivity
[…]
@@
from which it follows that $H=H^{\dagger}$, guaranteeing that the Lorentzian evolution $e^{iHt}$ preserves the norm. (More generally, if one assumes the Euclidean partition function is a covariant functional of the metric, one could deduce that Lorentzian evolution with respect to arbitrary lapse and shift functions will also be unitary.) Some constraints from only having states with positive norm were explored in [100, 101], and explained in a particularly clear way in [102].
@@ 3.1 Global de Sitter
where the $z^{i}$ describe a unit $S^{d-1}$ satisfying $z_{i}z^{i}=1$ and are expressible in terms of $d-1$ angular variables, $\Phi_{i}$.
@@
where $T\in\mathbb{R}$ and $\mathop{}\!\mathrm{d}\Omega_{d-1}^{2}$ is the round metric on the unit $d-1$-sphere. We can see from the metric that constant $T$ surfaces are $d$-spheres which shrink from the past asymptotic boundary $\mathcal{I}^{-}$ at $T=-\infty$ to $T=0$ and grow from $T=0$ to the future asymptotic boundary $\mathcal{I}^{+}$ at $T=+\infty$. The embedding space and global de Sitter coordinates are related as

### v3

@@ 1 Introduction
The reason we can obtain these stronger constraints, is that we are using the additional conditions that (i) we are considering the vacuum state only, and (ii) the dynamics of the theory come from a local Lagrangian that has $\bf{CRT}$ symmetry around every point. This combination of discrete symmetries with locality turns out to give results much more powerful than those which can be obtained by simply thinking of discrete symmetries as acting globally on correlators, without considering the aetiology of the correlators, i.e. the manner in which they arise from interactions, starting from a Bunch-Davies boundary condition.
@@ 1.2 Applications to Cosmology
$\text{$\bf{RR}$}:$ $\left[\psi^{(L)}_{n}(\eta_{0};\textbf{k})\right]^{*}\!$ $=\,\,e^{i\pi\left((d+1)L-1\right)}\psi^{(L)}_{n}(\eta_{0};e^{-i\pi}\textbf{k})\,.$ (1.9)
@@
where the last term indicates that $\bf{CRT}$ can only fix the phase up to an overall $\pm$ sign. $\Delta$ is the conformal dimension of the $n$ operators $\overline{\phi}_{+}$ used to differentiate the boundary wavefunction $\overline{\Psi}[\overline{\phi}_{-}]$, which is written in the basis of the conjugate operators $\overline{\phi}_{-}$ with dimension $d-\Delta$. [^9] From a dS/CFT perspective, the objects with dimension $d-\Delta$ are the “sources” of the CFT partition function $Z_{\text{CFT}}=\overline{\Psi}$, while the ones with dimension $\Delta$ are the “operators”. But from a cosmology perspective, both are operators in the bulk Hilbert space, and we have to choose a basis. In cosmology, one normally chooses this basis by writing $\overline{\Psi}$ as a function of whichever field component falls off more slowly, which would imply that $\Delta\geq d/2$, but the phase formula above would still be valid if we make the opposite choice. We do not attempt in this paper to extend our results to heavy fields in the principal series, where $\Delta=d/2+i\nu$. Doing so would require making the unpleasant decision to either sacrifice self-adjointness, or conformal invariance of the boundary conditions. See e.g. [76, 77, 78, 79, 80, 81, 82, 83, 84, 85] for discussions regarding the principal series. This result holds for any boundary wavefunction coefficients that ae UV- and IR-finite (i.e. free from logarithmic divergences in $\eta$ or the UV cutoff $\epsilon$), as well as the coefficient in front of the leading $\log(\eta)$ divergence for any IR-divergent wavefunction coefficients, or the leading $\log(\epsilon)$ divergence for any UV-divergent diagram. [^10] There is also an extra correction for diagrams with spinor propagators, see footnote 43.
@@
This then tells us that massless scalar fields and gravitons have real boundary wavefunction coefficients to all loop order in even spacetime dimension provided they don’t diverge in the UV. When UV divergences are present one finds in dimensional regularisation (dim-reg) a generic complex phase proportional to the dim-reg parameter $\delta$, i.e. $\arg(\overline{\psi}^{(L)}_{n})\propto\pi\delta$. For tree-level, IR-finite and scale-invariant Feynman-Witten diagrams in cosmology, this was recently found independently [74, 75] where the authors proved this reality condition by direct analysis of the bulk time integrals. [^11] AT would like to express his gratitude to David Stefanyszyn, Xi Tong and Yuhang Zhu for various extensive and helpful discussions on reality in cosmology. The implications of this for cosmology in terms of no-go theorems for parity violations has previously been explored for the specific case of the tree-level scalar trispectrum in [86, 87], and has been treated for more general tree-level and loop diagrams in [88]. This is also consistent with what is found in flat space amplitudes, which can be shown to be real at tree-level, see e.g. [89]. [^12] AT thanks Carlos Duaso Pueyo for making him aware of this result in amplitudes.
@@ 1.3 Implications for dS/CFT
[…]
@@
  • For $d=2$ at tree-level, $c=i\frac{3\ell}{2G_{N}}$ (e.g. from applying Friedel’s analysis to dS [59]); while for the case of pure gravity at 1-loop order we have $c=i\frac{3\ell}{2G_{N}}+13+\mathcal{O}(G_{N})$ [^15] The renormalisation of Newton’s constant $G_{N}$ is scheme dependent, meaning its exact running or loop corrections depend on the choice of renormalisation procedure. In (A)dS3, when one reformulates the gravitational path integral (GPI) using a boundary Schwarzian-type description, there exists a localisation result (analogous to the Duistermaat–Heckman theorem) which implies that, in that specific scheme, the renormalisation of $G_{N}$ stops at one loop. However, this one-loop exactness is not universal and might not hold in other renormalisation schemes. Importantly, the physical quantity is the boundary central charge (which appears in the dual CFT), not the bare $G_{N}$ itself. Therefore, as is standard in QFT, one should treat $G_{N}$ as a scheme-dependent parameter that is adjusted (renormalised) to keep physical observables—like the central charge—fixed. [91]; [^16] We thank Jordan Cotler and Kristan Jensen for making us aware of their previous result for the 1-loop contribution to the central charge in $dS_{3}$ as well as upcoming work for loop calculations in $dS_{2}$ and $dS_{4}$. The more observant reader will notice that the real $1$-loop contribution is positive in the convention chosen by the authors as predicted from the combination of the normalisability constraint and the phase formula.
@@ 2.2 Outline of CPT Theorem
Assumption 3 allows us to analytically continue to Euclidean signature, using the fact that $e^{-H\Delta\tau}$ is suppressed for positive Euclidean times $\Delta\tau>0$. Let us adopt coordinates in Euclidean signature where $\tau=it$ and $x_{I}$, I=1dI=1\,...\,d are the spatial coordinates. We highlight here (as it will become important in the cosmological case) that the $\bf{R}$ in $\bf{CRT}$ reflects a single spatial coordinate, say $x_{1}\rightarrow-x_{1}$.
The $SO^{+}(d,1)$ Lorentz Invariance in $d+1$-dimensional Minkowski spacetime analytically continues to $SO(d+1)$ rotational invariance in Euclidean signature. On the other hand, Unitarity implies that Euclidean correlators satisfy Reflection Positivity, which we will define in the following section. However, it is not necessary to enforce either of these relationships in their entirety to ensure $\bf{CRT}$ symmetry. We do not require full $SO^{+}(d,1)$ Lorentz Invariance; we simply need the subgroup $SO^{+}(1,1)$ boost  [^18] Consider a Lorentz boost in the $t$-$x^{d}$ plane of $d+1$-dimensional Minkowski spacetime. The infinitesimal transformation of coordinates is given by $\begin{pmatrix}t^{\prime}\\ x^{D}{}^{\prime}\end{pmatrix}=\begin{pmatrix}\cosh\xi&-\sinh\xi\\ -\sinh\xi&\cosh\xi\end{pmatrix}\begin{pmatrix}t\\ x^{D}\end{pmatrix}$, (2.1) where $\xi$ is the rapidity parameter of the boost, and we have omitted dependence on the other d1d-1 spatial coordinates for simplicity. To Wick rotate to Euclidean signature, we set $t=-i\tau\qquad\text{and}\qquad\xi=-i\theta\,$, (2.2) which leads to the transformation $\begin{pmatrix}-i\tau^{\prime}\\ x^{D}{}^{\prime}\end{pmatrix}=\begin{pmatrix}\cosh(-i\theta)&\sinh(-i\theta)\\ \sinh(-i\theta)&\cosh(-i\theta)\end{pmatrix}\begin{pmatrix}-i\tau\\ x^{D}\end{pmatrix}$. (2.3) Using the identities $\cosh(-i\theta)=\cos\theta$ and $\sinh(-i\theta)=-i\sin\theta$, this becomes $\begin{pmatrix}\tau^{\prime}\\ x^{D}{}^{\prime}\end{pmatrix}=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\begin{pmatrix}\tau\\ x^{D}\end{pmatrix}$, (2.4) which is a clockwise rotation in the $\tau$-$x^{D}$ plane by angle $\theta$. Hence, under the Wick rotation $t=-i\tau$ and $\xi=-i\theta$, the non-compact Lorentz boosts $SO^{+}(1,1)$ in Minkowski space analytically continue to compact Euclidean rotations $SO(2)$ in the $\tau$-$x^{D}$ plane. (this will become important when studying the implications of the CPT theorem for cosmology). In fact, all that is required is a discrete, $\mathbb{Z}_{2}$ symmetry from each of them:
@@
  • A $180^{\circ}$ ($\pi$) rotation in the $\tau$-$x$ plane corresponds to the transformation: [^20] Some authors refer to this transformation as “CPT”. The difference in convention goes back to the earliest papers on the CPT theorem, see [95] for a nice review. In the language of [95], we are defining $\bf{T}$ as a “Wigner time reversal” (which involves a complex conjugation) while other authors use “Schwinger time reversal” (which instead reverses the order of operators. For example, the recent paper by Susskind [96] and [97], as well as Harlow and Numasawa [98] on gauging “CRT”, is actually defining $\bf{CRT}$ as what we call the $180^{\circ}$ rotation, at least in their Euclidean signature construction (see also [99], where the authors find that in order to construct physically consistent bulk local states, it is necessary to take a CPT-invariant linear combination of the two primary states of the dual CFT). It is unclear whether it makes sense to gauge our Wignerian $\bf{CRT}$, as this would raise the question of what it means to have a non-trivial complex conjugation of amplitudes when going around a non-contractible cycle. (Perhaps it just means the whole amplitude vanishes.)
@@
For theories involving spinor fields, it is also necessary to consider projective representations of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. In this paragraph, let us assume for simplicity that the theory is fully rotationally covariant. Then a $360^{\circ}$ rotation gives a factor of $(-1)^{2s}$ (where $s$ is the spin); this is of course the same as the fermion parity $(-1)^{F}$ for fields that satisfy spin-statistics. Hence, the $180^{\circ}$ rotation $R_{\pi}$ has order $4$ when acting on half-integer spin fields, whilst Unitarity requires that ${\bf RR}^{2}=+1$ and hence in a rotationally covariant theory we also need ${\bf{CRT}}^{2}=+1$. This is only possible if the group anticommutes in the sense that $A,B]:=ABA^{-1}B^{-1}=-1$ when $A,B$ are any two distinct elements of $\bf{RR}$, $\bf{CRT}$, ${\bf R}_{\pi}$}. One thus finds that the appropriate double cover of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ for spinors is the dihedral group $D_{4}$. [^22] With the addition of one additional magical fact that Unitarity requires ${\bf CRT}^{2}=(-1)^{2s}(-1)^{F}$, which requires a close inspection of the transformation properties of 2-pt correlators to prove, the considerations in this paragraph provide all the necessary ingredients to prove the spin-statistics theorem. But spin-statistics requires at least covariance under $90^{\circ}$ rotations, not just $180^{\circ}$ rotations, so it is less robust than the CPT theorem. We will focus on fields with integer spin representations and not consider half-integer spin fields in detail this paper, but a similar formalism should also provide a much easier approach to obtain a Unitarity constraint for correlators involving spinors in flat space.
@@ 2.3 Reflection Positivity
For a fixed number, $N$, of insertions on each side, we would obtain the positivity of some $2N$-point function. However, it is also allowed to consider states coming from quantum superpositions of different values of $N$. Because the sums are taken independently on both sides of the $\tau=0$ surface, Reflection Positivity also places constraints on the correlations with odd numbers of points. It is not, however, allowed for the operators which define the ket to stray “offside” onto the wrong side of the $\tau=0$ surface, even if you include the corresponding reflected operator.
@@
from which it follows that $H=H^{\dagger}$, guaranteeing that the Lorentzian evolution $e^{iHt}$ preserves the norm. (More generally, if one assumes the Euclidean partition function is a covariant functional of the metric, one could deduce that Lorentzian evolution with respect to arbitrary lapse and shift functions will also be unitary.) Some constraints from only having states with positive norm were explored in [100, 101, 102, 103, 104, 105, 106], and explained in a particularly clear way in [107].
@@ 3.1 Global de Sitter
where the $z^{i}$ describe a unit $S^{d-1}$ satisfying $z_{i}z^{i}=1$ and are expressible in terms of d1d-1 angular variables, $\Phi_{i}$.
@@
where $T\in\mathbb{R}$ and $\mathop{}\!\mathrm{d}\Omega_{d-1}^{2}$ is the round metric on the unit d1d-1-sphere. We can see from the metric that constant $T$ surfaces are $d$-spheres which shrink from the past asymptotic boundary $\mathcal{I}^{-}$ at $T=$-$\infty$ to $T=0$ and grow from $T=0$ to the future asymptotic boundary $\mathcal{I}^{+}$ at $T=+\infty$. The embedding space and global de Sitter coordinates are related as